OpenStudy (anonymous):

Greatest common factor 54, 45 40,10 68,34 55,75 25,55 54,27

OpenStudy (anonymous):

SOME ONE HELP ME PLEASE(:

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

I'll do the first one to get you started We can factor 54 into 9*6 then we can factor 9 into 3*3 then we can factor 6 into 2*3 So 54 = 9*6 = (3*3)*(2*3) = 2*3*3*3 This means that 54 = 2*3*3*3 Use the same technique to factor 45 45 = 3*3*5 So we have 54 = 2*3*3*3 45 = 3*3*5 Now highlight/circle all the terms that are shared between both numbers. There are two 3's shared So the GCF of 54 and 45 is 3*3 = 9

OpenStudy (anonymous):

What? ^ rahul1995 I'm confused

OpenStudy (anonymous):

ok just divide both numbers together till they have common factors. I divided 40 and 10 both by 2 first, getting 20 and 5, now divide both by 5, get 4 and 1. can't divide the further. So 2*5=10 is our answer. It's called long division method (sorry i put 10 instead of 4 in the drawing xD)

OpenStudy (anonymous):

this one is correct one